OpenStudy (anonymous):

how do i find dot-product and angle between v and w in this? v=3i-j w=i+j

jimthompson5910 (jim_thompson5910):

if v = ai + bj and w = ci + dj then v dot w = a*c + b*d

OpenStudy (anonymous):

Angle between them = \[\cos^{-1} \frac{ d.product }{ |v|*|w| }\]

OpenStudy (anonymous):

thanks, you where both very helpful. :)